# Infinity-Parser: Layout-Aware Reinforcement Learning for Scanned Document Parsing

### v1

@@ 3.2 Main Results
Table 6: End-to-end text recognition performance on OmniDocBench: evaluation using edit distance across 9 PDF page types. We compare with Mathpix, MinerU [52], Marker [37], GOT-OCR [59], Nougat [6], GPT-4o [1], Qwen2-VL-72B [55], InternVL2-76B [8], Qwen2.5-VL-7B [4], InternVL3-8B [75].
Diverse Page Types Evaluation To further investigate model behavior across diverse document types, we evaluated text recognition performance on nine distinct page categories. As shown in Table 6, pipeline-based systems such as MinerU [52] and Mathpix achieved strong results on structured formats like academic papers and financial reports. General-purpose vision-language models (VLMs) demonstrated better generalization on less formal page types, including presentation slides and handwritten notes. However, for challenging formats such as newspapers, most VLMs underperformed, while pipeline tools maintained relatively lower error rates. Notably, our proposed Infinity-Parser-7B achieved consistently low edit distances across all document types, outperforming both pipeline-based systems and general-purpose VLMs in overall accuracy. This highlights the robustness and adaptability of our reinforcement learning approach across diverse and complex document layouts.
Model Anchor ArXiv Math Tables Scans H&F MultiCol TinyText Base Overall
GOT OCR 52.7 52.0 0.2 22.1 93.6 42.0 29.9 94.0 48.3
Marker v1.6.2 24.3 22.1 69.8 24.3 87.1 71.0 76.9 99.5 59.4
MinerU v1.3.10 75.4 47.4 60.9 17.3 96.6 59.0 39.1 96.6 61.5
Mistral OCR API 77.2 67.5 60.6 29.3 93.6 71.3 77.1 99.4 72.0
GPT-4o 51.5 75.5 69.1 40.9 94.2 68.9 54.1 96.7 68.9
GPT-4o 53.5 74.5 70.0 40.7 93.8 69.3 60.6 96.8 69.9
Gemini Flash 2 32.1 56.3 61.4 27.8 48.0 58.7 84.4 94.0 57.8
Gemini Flash 2 54.5 56.1 72.1 34.2 64.7 61.5 71.5 95.6 63.8
Qwen 2 VL 19.7 31.7 24.2 17.1 88.9 8.3 6.8 55.5 31.5
Qwen 2.5 VL 63.1 65.7 67.3 38.6 73.6 68.3 49.1 98.3 65.5
olmOCR v0.1.68 72.1 74.7 71.5 43.7 91.6 78.5 80.5 98.1 76.3
olmOCR v0.1.68 75.6 75.1 70.2 44.5 93.4 79.4 81.7 99.0 77.4
[…]
Table 7: Performance comparison on the olmOCR [39] benchmark across diverse document structures. ArXiv: scientific articles. Math: scanned mathematical content. Scans: historical scanned documents. H&F: headers and footers. MultiCol: multi-column layout. TinyText: long lines or extremely small text. Base: base text recognition accuracy. Higher is better. The Anchor column indicates whether external PDF metadata (e.g., layout hints) is used to assist layout prediction.
Document-Level OCR Evaluation on olmOCR-Bench Table 7 reports performance on the olmOCR-Bench benchmark, which evaluates document-level OCR across diverse layouts and domains. Infinity-Parser-7B achieves the highest overall score (79.1), followed closely by olmOCR v0.1.68 (Anchored) (77.4), both demonstrating strong performance in complex categories like multi-column layouts and scanned math content. The results highlight the effectiveness of anchored prompting, with anchored versions of models (e.g., GPT-4o, olmOCR) significantly outperforming their non-anchored counterparts—especially on tables and old scans. This underscores the importance of layout-aware extraction techniques. In contrast, traditional pipelines like Marker and GOT OCR lag behind in structural accuracy, reinforcing the value of modern VLM-based approaches in high-fidelity PDF understanding.
3.3 Ablation Study
We perform ablation experiments to evaluate the individual contributions of our three core design choices: (1) data quality verification and (2) multi-aspect rewards. We report all ablation results using two primary evaluation metrics: OverallEdit and OverallCat.. OverallEdit represents the average edit-based overall score across English and Chinese pages, as shown in Table 2. In contrast, OverallCat. reflects the mean category-level performance across nine types of scanned documents pages, following the same evaluation setting as Table 10 in the Appendix. To save computational resources, the ablation studies are conducted on a randomly selected subset of 22K samples, rather than using the full 55K dataset as in the final results.
Effect of Data Validity Table 8 highlights two important findings regarding data effectiveness and the limitations of the SFT paradigm. First, increasing the size of training data improves performance in terms of OverallCat., which drops significantly from 0.257 at 5k to 0.166 at 22k, confirming the value of larger datasets in enhancing content categorization accuracy. However, the OverallEdit metric shows a non-monotonic trend: while initial gains are observed from 5k to 10k, performance begins to degrade as the dataset size increases further. This suggests that beyond a certain point, the SFT paradigm struggles to maintain structural precision, possibly due to the lack of layout-aware supervision, revealing its limitations in scaling for layout-sensitive document parsing.
As shown in Table 9, we evaluate the impact of data scale on RL training with layout-aware rewards. The zero-shot baseline performs worst, highlighting the importance of training data. From 5k to 15k, category error steadily decreases and edit distance improves, indicating better generalization. At 22k, performance peaks, though minor edit distance fluctuations suggest possible noise. These results show that scaling data enhances RL’s structural modeling capabilities.
Method Data Size OverallEdit $\downarrow$ OverallCat. $\downarrow$
Zero Shot 0.346 0.259
SFT 5k 0.332 0.257
SFT 10k 0.353 0.167
SFT 15k 0.374 0.179
SFT 22k 0.375 0.166
Table 8: Results with different data sizes.
Method Data Size OverallEdit $\downarrow$ OverallCat. $\downarrow$
Zero-Shot - 0.346 0.259
RL 5k 0.275 0.168
RL 10k 0.310 0.153
RL 15k 0.324 0.147
RL 22k 0.260 0.123
Table 9: Results with different data types.
Effect of Data Construction Methods Table 10 demonstrates that synthetic data serves as an effective complement to real-world data for document parsing. Under both SFT and RL settings, synthetic data consistently achieves lower scores in OverallEdit and OverallCat., indicating better structural alignment and parsing accuracy. Notably, combining real and synthetic data under the RL setup yields the best overall performance,
Method Data Type OverallEdit $\downarrow$ OverallCat. $\downarrow$
SFT Real 0.337 0.279
SFT Synth 0.290 0.214
RL Real 0.285 0.164
RL Synth 0.262 0.156
RL Synth+Real 0.273 0.155
Table 10: Results with different data sizes.
especially in OverallCat. (0.155), suggesting that the structural precision of synthetic documents enhances the diversity and complexity offered by real-world samples. These results highlight the importance of leveraging both data sources to achieve robust and layout-aware document understanding.
Method Edit Dist. Reward Count Reward Order Reward SFT RL OverallEdit $\downarrow$ OverallCat. $\downarrow$
Zero Shot - - - - - 0.346 0.259
SFT - - - 22k - 0.375 0.166
SFT + RL - - 5k 17k 0.345 0.244
Zero + RL - - - 22k 0.287 0.156
Zero + RL - - 22k 0.280 0.141
Zero + RL - 22k 0.260 0.123
Table 11: Results under different reward designs.
Effect of Multi-Aspect Rewards Table 11 shows that reinforcement learning consistently outperforms supervised fine-tuning, particularly under layout-aware reward designs. Compared to the SFT baseline (0.375 / 0.166), the RL method with distance-based reward (Zero + $R_{\text{dist}}$) achieves better OverallEdit (0.287) and comparable OverallCat. (0.156). Further introducing count and order rewards leads to significant improvements in structural consistency, with Zero + $R_{\text{dist}}+R_{\text{count}}$ achieving 0.280 / 0.141, and the full reward setting reaching 0.260 / 0.123. These results highlight RL’s ability to incorporate structural inductive biases via reward design, enabling better alignment with task-specific goals and reducing overfitting to token-level patterns. The improvements across reward configurations confirm the effectiveness of multi-aspect rewards in enhancing layout-aware generalization.
@@ 3.3 Ablation Study
(a) Comparison of four Markdown generation results on a single case, illustrating progressive improvement from direct inference to full reward integration.
(b) Markdown extraction from a PowerPoint-style PDF slide. The figure includes the original slide (a), human-annotated ground truth (b), and outputs from MinerU, GPT-4o, and Infinity-Parser (c–e). MinerU and GPT-4o introduce spacing and formatting errors. In contrast, Infinity-Parser accurately preserves dialogue structure and layout fidelity.
@@
(a) Comparison of Markdown extraction from academic literature. The figure shows the original PDF (a), ground truth annotations (b), and extraction results from MinerU, GPT-4o, and Infinity-Parser (c–e). Infinity-Parser produces the most accurate output.
(b) Comparison of Markdown extraction from a book-style PDF. GPT-4o introduces multiple errors; MinerU retains unnecessary line breaks; Infinity-Parser produces clean and accurate output.
(a) Markdown extraction from a colorful textbook-style PDF. The figure displays the original illustrated page (a), ground truth annotations (b), and outputs from MinerU, GPT-4o, and Infinity-Parser (c–e). MinerU and GPT-4o introduce issues such as redundant content, missed spacing, and incorrect formatting. In contrast, Infinity-Parser accurately identifies section titles and body text, preserving the structure and semantics of the original content.
@@
(b) Markdown extraction from an exam-style PDF. The figure presents the original exam document (a), ground truth annotations (b), and the extraction results from MinerU, GPT-4o, and Infinity-Parser (c–e). GPT-4o and MinerU both introduce redundant text and formatting errors, such as incorrect title levels and misplaced content. In contrast, Infinity-Parser accurately captures the heading hierarchy and structured list format.
@@
(a) Markdown extraction from a magazine-style PDF. The figure shows the original visually-rich page (a), ground truth annotations (b), and results from MinerU, GPT-4o, and Infinity-Parser (c–e). GPT-4o suffers from significant formatting errors, while Infinity-Parser accurately preserves the structural hierarchy and formatting.
@@
(b) Markdown extraction from a newspaper-style PDF. This figure presents the original densely formatted page (a), ground truth annotations (b), and the results from MinerU, GPT-4o, and Infinity-Parser (c–e). Both MinerU and GPT-4o exhibit issues such as incorrect title levels, redundant content, and format errors. In contrast, Infinity-Parser accurately identifies the main title, maintains structural formatting, and preserves content integrity.
3.4 Result Analysis
Figure 5(a) illustrates a progressive improvement in Markdown generation quality across different training strategies. The zero-shot model fails to capture key structural elements, omitting titles and producing redundant or incomplete content. With SFT, the model better identifies section headers and general layout but still suffers from symbol-level errors and repeated outputs. In contrast, the layout-aware RL model demonstrates the most accurate and coherent result, successfully preserving the document hierarchy and eliminating redundancy. This highlights the effectiveness of layout-aware rewards in guiding the model toward semantically and structurally faithful document parsing. Infinity-Parser consistently outperforms baselines across a wide range of document types, including academic papers, books, textbooks, exams, magazines, government notices, newspapers, and slides. It excels in structure parsing, content recognition, and formatting accuracy, demonstrating strong generalization and layout robustness (Figures 5(b) - 8(b)).
4 Related Work
4.1 Reinforcement Learning for Language Models
Recent advancements in Large Language Models (LLMs) such as OpenAI’s GPT series [35]), DeepSeek-R1 [11], and Gemini [47] have highlighted the significant potential of Reinforcement Learning (RL) in enhancing their reasoning capabilities. This RL paradigm has been successfully extended to other domains demanding sophisticated reasoning, including code generation [23, 68], autonomous tool utilization [41, 53], and information retrieval [33]. Similarly, RL has demonstrated its efficacy in the domain of Visual Language Models (VLMs), including precise object counting [38], nuanced visual perception [29], and complex multimodal reasoning (e.g., Pixel Reasoner [46], VL-Rethinker [54], Vision-R1 [16]). These pioneering works have predominantly relied on binary outcome rewards to guide RL training. Complementary to these efforts, our work demonstrates the effectiveness of incorporating layout-aware and layout-based rewards for document parsing, offering a more granular and contextually relevant feedback mechanism.
4.2 VLM-based Document Parsing
Recent advancements in document understanding and optical character recognition (OCR) have highlighted their importance as critical benchmarks for evaluating the perceptual capabilities of vision-language models (VLMs). By incorporating large-scale OCR corpora during pretraining, models such as GPT-4o [1] and Qwen2-VL [2] have achieved competitive performance on document content extraction tasks. Building upon these foundations, the emergence of VLMs has further accelerated the progress of end-to-end document parsing, giving rise to a range of models such as Donut [6], Nougat [5], Kosmos-2.5 [31], Vary [61], mPLUG-DocOwl [15], Fox [25], and GOT [59]. These models have continued to improve their understanding of visual layouts and textual content by leveraging advancements in visual encoders [10], language decoders [2], and data construction pipelines. Despite the success of these VLM-based approaches in enabling end-to-end document parsing, they still face generalization challenges on downstream layout parsing tasks [52, 59]. To address this issue, we propose leveraging reinforcement learning to provide a more effective training paradigm that better aligns with the demands of document parsing.
5 Conclusion
In this work, we presented layoutRL, the first end-to-end reinforcement-learning framework that makes document parsers explicitly layout-aware by optimizing a multi-aspect reward—combining normalized edit-distance, paragraph-count accuracy, and reading-order preservation. To support this training paradigm, we released Infinity-Doc-55K, a 55K document corpus blending high-fidelity synthetic scanned document parsing data with expert-filtered real-world samples, providing the large-scale, precise supervision end-to-end models require. We implement the proposed method as a vision-language-model–based parser, Infinity-Parser. Powered by layoutRL, it achieves new state-of-the-art performance among end-to-end models on English and Chinese tasks, including OCR, table and formula extraction, and reading order detection. By releasing our code, dataset, and trained model, we hope to catalyze further advances in robust, reliable document understanding.

### v3

@@ Appendix E More Results
Table 9: End-to-end text recognition performance on OmniDocBench: evaluation using edit distance across 9 PDF page types. We compare with Mathpix, MinerU (Wang et al., 2024b), Marker (Paruchuri, 2024), GOT-OCR (Wei et al., 2024), Nougat (Blecher et al., 2024), GPT-4o (Brown et al., 2020), Qwen2-VL-72B (Wang et al., 2024c), InternVL2-76B (Chen et al., 2024), Qwen2.5-VL-7B (Bai et al., 2025), InternVL3-8B (Zhu et al., 2025).
Model Language Table Frame Type Special Situation Overall $\uparrow$
EN ZH Mixed Full Omission Three Zero Merge Cell(+/-) Formula(+/-) Colorful (+/-) Rotate(+/-)
PaddleOCR (Li et al., 2022a) 76.8 71.8 80.1 67.9 74.3 81.1 74.5 70.6/75.2 71.3/74.1 72.7/74.0 23.3/74.6 73.6
RapidTable (RapidAI, 2023) 80.0 83.2 91.2 83.0 79.7 83.4 78.4 77.1/85.4 76.7/83.9 77.6/84.9 25.2/83.7 82.5
StructEqTable (Zhou et al., 2024) 72.8 75.9 83.4 72.9 76.2 76.9 88.0 64.5/81.0 69.2/76.6 72.8/76.4 30.5/76.2 75.8
GOT-OCR (Wei et al., 2024) 72.2 75.5 85.4 73.1 72.7 78.2 75.7 65.0/80.2 64.3/77.3 70.8/76.9 8.5/76.3 74.9
Qwen2-VL-7B (Wang et al., 2024c) 70.2 70.7 82.4 70.2 62.8 74.5 80.3 60.8/76.5 63.8/72.6 71.4/70.8 20.0/72.1 71.0
InternVL2-8B (Chen et al., 2024) 70.9 71.5 77.4 69.5 69.2 74.8 75.8 58.7/78.4 62.4/73.6 68.2/73.1 20.4/72.6 71.5
Qwen2.5-VL-7B (Wang et al., 2024c) 87.4 80.7 93.5 86.4 85.1 84.1 88.7 77.5/89.8 82.1/87.2 77.1/87.5 56.5/86.0 85.5
InternVL3-8B (Zhu et al., 2025) 79.5 86.0 91.7 85.5 80.7 83.9 85.9 71.9/90.9 74.0/86.7 82.1/85.3 12.6/85.5 84.3
Infinity-Parser-7B 84.7 86.7 94.8 85.5 86.5 87.4 89.4 78.6/90.7 81.9/ 87.5 83.2/88.0 68.8/86.7 86.4
Appendix F Prompt Strategy for Parsing Tasks.
Prompt Template summarizes the prompt designs for two key parsing tasks: document parsing and table parsing. For document parsing, the prompts instruct the model to recognize visual regions and convert their contents into structured Markdown. This design ensures consistent region-level extraction across documents with diverse layouts.
For table parsing, although the prompts are phrased differently, they share the same objective: transforming table content from images into HTML. This diversity encourages the model to generalize across variations in phrasing and reduces overfitting to a single instruction template. Notably, HTML is used here to match the evaluation format, but the resulting outputs can be easily converted to Markdown if needed for downstream use.
Appendix G Case Analysis
Figure 8 illustrates a progressive improvement in Markdown generation quality across different training strategies. The zero-shot model fails to capture key structural elements, omitting titles and producing redundant or incomplete content. With SFT, the model better identifies section headers and general layout but still suffers from symbol-level errors and repeated outputs. In contrast, the layout-aware RL model demonstrates the most accurate and coherent result, successfully preserving the document hierarchy and eliminating redundancy. This highlights the effectiveness of layout-aware rewards in guiding the model toward semantically and structurally faithful document parsing.
@@ Appendix G Case Analysis
Figure 8: Comparison of four Markdown generation results on a single case, illustrating progressive improvement from direct inference to full reward integration.
Infinity-Parser exhibits consistent improvements across a wide spectrum of document types, including academic papers, books, colorful textbooks, exam papers, magazines, government notices, newspaper articles, and PowerPoint-style slides. These gains are reflected in structural parsing, title and content recognition, formatting accuracy, and robustness to diverse visual layouts. As shown in Figures 9 through 14, we provide detailed visual comparisons with existing models, where Infinity-Parser consistently achieves superior results. These findings underscore the effectiveness and generalizability of our layout-aware RL approach across complex, real-world PDF formats.
@@
Figure 9: Comparison of Markdown extraction from academic literature using different models. The figure shows the original PDF (a), ground truth annotations (b), and extraction results from three models: MinerU, GPT-4o, and Infinity-Parser (c–e). Infinity-Parser produces the most accurate output, correctly identifying titles and content while avoiding redundancy and omissions.
Figure 10: Comparison of Markdown extraction from a book-style PDF using different models. This figure displays the original document (a), human-annotated ground truth (b), and the extraction results from MinerU, GPT-4o, and Infinity-Parser (c–e). GPT-4o introduces multiple errors, such as incorrect headings, redundancy, and missing lines. MinerU retains unnecessary line breaks and repeated text. In contrast, Infinity-Parser correctly identifies the content structure, including titles and paragraphs, producing clean and accurate output.
@@
Figure 11: Comparison of Markdown extraction from an exam-style PDF. The figure presents the original exam document (a), ground truth annotations (b), and the extraction results from MinerU, GPT-4o, and Infinity-Parser (c–e). GPT-4o and MinerU both introduce redundant text and formatting errors, such as incorrect title levels and misplaced content. In contrast, Infinity-Parser accurately captures the heading hierarchy and structured list format, faithfully reproducing the content as intended in the original layout.
@@
Figure 12: Comparison of Markdown extraction from a magazine-style PDF. The figure shows the original visually-rich page (a), ground truth annotations (b), and results from MinerU, GPT-4o, and Infinity-Parser (c–e). Due to the complex layout and dark background, GPT-4o suffers from significant formatting errors. Infinity-Parser accurately preserves the structural hierarchy and formatting, demonstrating robustness in handling stylized layouts.
@@
Figure 13: Comparison of Markdown extraction from a newspaper-style PDF. This figure presents the original densely formatted page (a), ground truth annotations (b), and the results from MinerU, GPT-4o, and Infinity-Parser (c–e). Due to the complex multi-column layout and title hierarchy, both MinerU and GPT-4o exhibit issues such as incorrect title levels, redundant content, and format errors. In contrast, Infinity-Parser accurately identifies the main title, maintains structural formatting, and preserves content integrity, demonstrating strong layout understanding in challenging document types.
Figure 14: Comparison of Markdown extraction from a PowerPoint-style PDF slide. The figure includes the original slide (a), human-annotated ground truth (b), and outputs from MinerU, GPT-4o, and Infinity-Parser (c–e). MinerU and GPT-4o both struggle with layout fidelity, introducing spacing and formatting errors. In contrast, Infinity-Parser preserves the dialogue structure and formatting accurately, demonstrating its capability to handle informal, visually decorated slides effectively.